how do I solve?
3*5^(2t-1)=75

Answer by Jk22(389)

3*5^(2t-1) = 75 | dividing by 3

5^(2t-1) = 25 = 5^2 | equating exponents, or log()/log(5)

2t - 1 = 2, | adding 1

2t = 3 |dividing by 2

t = 3/2